OpenStudy (anonymous):

simplify (8x) -2 to the second power

OpenStudy (luigi0210):

\[a^{-1}=\frac{1}{a}\]

OpenStudy (anonymous):

Im still confused with you answer and my question

OpenStudy (mertsj):

The - part of the exponent means "the reciprocal of"

OpenStudy (mertsj):

So (8x)^-2 means "the reciprocal of 8x" raised to the second power.

OpenStudy (anonymous):

ok so would that mean the answer is 8*8?

OpenStudy (mertsj):

What is the reciprocal of 8x?

OpenStudy (anonymous):

the problem just asks simplify (8x)^-2

OpenStudy (mertsj):

What is the reciprocal of 8x?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

I am aware of what the problem says.

OpenStudy (anonymous):

ok would it be -64

OpenStudy (mertsj):

What is the reciprocal of 8x?

OpenStudy (anonymous):

1

OpenStudy (mertsj):

So the reason you are having trouble is that you don't know what a reciprocal is?

OpenStudy (anonymous):

correct

OpenStudy (mertsj):

Let me show you some examples:

OpenStudy (anonymous):

ok great

OpenStudy (mertsj):



OpenStudy (mertsj):

So, I ask you for the fourth and final time....what is the reciprocal of 8x ?

OpenStudy (anonymous):

1/8x

OpenStudy (mertsj):

Now we are making progress.

OpenStudy (mertsj):

\[(8x)^{-2}\] means take the reciprocal of 8x and square it.

OpenStudy (anonymous):

ok im getting a better understanding now

OpenStudy (mertsj):

So now that we know what the reciprocal is, we have: \[(8x)^{-2}=(\frac{1}{8x})^2=\frac{1^2}{8^2x^2}=\frac{1}{64x^2}\]

OpenStudy (anonymous):

so instead of being 1/64^-2 it will be 1/64^2

OpenStudy (mertsj):

It will be: \[\frac{1}{64x^2}\]

OpenStudy (anonymous):

great thanks

OpenStudy (mertsj):

yw